A ball is thrown in the air from the top of a building. Its height, in meters above the ground, as a function of time in seconds, is given by
h(t) = −4.9t² + 24t + 8
How long does it take to reach the maximum height?
I'd appreciate a detailed written version of this problem, It's a bit confusing for me.

Found 2 solutions by ikleyn, mccravyedwin:
Answer by ikleyn(52794)

.
In this site, I prepared the lessons that cover the entire subject
    - Introductory lesson on a projectile thrown-shot-launched vertically up
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.
Read them attentively and learn how to solve this type of problems once and for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Projectiles launched/thrown and moving vertically up and dawn".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


DO  NOT  MISS  YOUR  CHANCE  (!)

It is really a  UNIQUE  OPPORTUNITY (!)



Answer by mccravyedwin(407)

When no time has passed, t for time is 0.  We substitute 0 for t 

h(t) = −4.9t² + 24t + 8
h(0) = −4.9(0)² + 24(0) + 8
h(0) = 0 + 0 + 8
h(0) = 8 meters

and find that the ball is 8 meters above the ground when it leaves the
pitcher's hand.

When 1 second has passed, we substitute 1 for t 

h(t) = −4.9t² + 24t + 8
h(1) = −4.9(1)² + 24(1) + 8
h(1) = -4.9(1)² + 24 + 8
h(1) = 27.1 meters

and find that the ball is 27.1 meters above the ground.

When 2 seconds have passed, we substitute 2 for t 

h(t) = −4.9t² + 24t + 8
h(2) = −4.9(2)² + 24(2) + 8
h(2) = -19.6 + 48 + 8
h(2) = 36.4 meters

and find that the ball is 36.4 meters above the ground.

When 3 seconds have passed, we substitute 3 for t 

h(t) = −4.9t² + 24t + 8
h(3) = −4.9(3)² + 24(3) + 8
h(3) = -44.1 + 72 + 8
h(3) = 35.9 meters

and find that the ball is 35.9 meters above the ground. So the ball is not as
high after 3 seconds as it was after 2 seconds.  That means it's already
reached its maximum height as is coming down.  So it must have reached its
maximum height at more than 2 seconds and less than 3 seconds.



To find that time we use the vertex formula to find the number of seconds that
must have passed when it reached its maximum height, -b/(2a) = -(24)/[2(-4.9)]
= 2.448979592 seconds.  That's the answer.  We didn't need to do all that work
at the top.  We could have just jumped to the vertex formula.  But you said you
wanted a detailed version, so you'd understand what was going on with the
height of the ball after 1 second, 2 seconds, and 3 seconds.

To find out how high it reached, we substitute t=2.448979592 seconds and get
that its maximum height was 37.3877551 meters.

Edwin